OpenStudy (anonymous):

Simplify -5n2+2(7m3+2n2)

OpenStudy (eli_moses):

\[-5n ^{2}+2(7m ^{3}+2n ^{2})\]

OpenStudy (eli_moses):

so from ..... \[-5n ^{2}+14m ^{3}+4n ^{2}\] now simplify that

OpenStudy (anonymous):

-1n2+14m3 ?

OpenStudy (camerondoherty):

Step 1 : Multiply 2 by 2n^2+7m^3 Factor: 2n^2+7m^3 A sum of two perfect cubes, a^3 + b^3 can be factored into : (a+b) • (a^2-ab+b^2) Checking : (a+b) • (a^2-ab+b^2) = a^3-a^2b+ab2+ba^2-b^2a+b^3 = a^3+(a^2b-ba^2)+(ab^2-b^2a)+b^3= a3+0+0+b3= a3+b3 Check : 2 is not a cube !! Ruling : Binomial can not be factored as the difference of two perfect cubes Equation at the end of step 1 : (0 - (5 • (n^2))) + 2 • (2n^2 + 7m^3) Step 2 : Simplify -5n^2 + 2•(2n^2+7m^3) factor as a Difference of Cubes: 2.1 Factoring: 14m^3-n^2 Theory : A difference of two perfect cubes, a^3 - b^3 can be factored into (a-b) • (a^2 +ab +b^2) Proof : (a-b)•(a^2+ab+b^2) = a^3+a^2b+ab^2-ba^2-b^2a-b^3 = a^3+(a^2b-ba^2)+(ab^2-b^2a)-b^3 = a^3+0+0+b^3 = a^3+b^3 Check : 14 is not a cube !! Ruling : Binomial can not be factored as the difference of two perfect cubes Answer: 14m^3 - n^2

OpenStudy (anonymous):

ok thanks :)

OpenStudy (camerondoherty):

sorry its long xD kinda overexagerated